What is the probability of drawing a 10 and then a 5 from a deck of cards (sampling WITH replacement)?

Answer by checkley77(12844)

DRAWING A 10
4/52=1/13
DRAWING A 5.
4/52=1/13
1/13*1/13=1/169=.0059=.59% ANS.